OpenStudy (anonymous):

solve each compound inqeuality and graph the solution set 5 less than or equal to 3a + 12 less than or equal to 16

OpenStudy (anonymous):

@abtster could you help again

OpenStudy (anonymous):

5 less than or equal to 3a + 12 less than or equal to 16 so 5 =< 3s + 12 and 3a + 12 =< 12 or so 5 =< 3s + 12 =< 12

OpenStudy (anonymous):

i got the first part:)

OpenStudy (anonymous):

the s is a typo, I meant an an the last is a 16 5 =< 3a + 12 =< 16 substract 12 from all parts -7 =< 3a =< 4 divide all parts by three - 7/3 =< a = < 4/3 that is - 2 1/3 =< a = < 1 1/3

OpenStudy (anonymous):

You still there @abtster ?

OpenStudy (anonymous):

Yes, i was trying to draw it

OpenStudy (anonymous):

ohh okay did you figure it out?

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

I don't know if the previous problem was answered right, all domains have the borders included so no open dots

OpenStudy (anonymous):